5) Monica grows coconuts and catches fish. Last year she harvested 1500 coconuts and 600 fish. She values one fish as havingSHOW WORK
Answer #1

Answer : 5) The answer is option E.

Based on given information,

3 coconuts = 1 fish

=> 1 coconut = 1 / 3 fish
=> 1500 coconut = (1/3) * 1500 = 500 fish

Monica gave 300 coconuts from her harvested coconuts of 1500 and 100 fish from her 600 fish to Rachel and Rachel consumes all. Monica stores 100 fish for next year harvest. So, total consumption of fish is (600 - 100) = 500 and the total consumption of coconut is 1500. As 1500 coconuts = 500 fish, hence total consumption in terms of fish is ( 500 + 500) = 1000 fish.

Therefore, option E is correct.